Your broker has offered to sell you some shares of Lucky Co. common stock that paid a dividend of $1 yesterday. You expect the dividend to grow at the rate of 2 percent per year for the next 3 years, and, if you buy the stock, you plan to hold it for 3 years and then sell it.

a. Find the expected dividend for each of the next three years.

b. If the appropriate discount rate is 8 percent, find the present value of the dividend stream for the next three years.

c. Suppose you expect to sell your stock three years from now at $30, what is the present value of $30? Assume the discount rate is 8 percent.

d. Find the most price you should pay for this stock at present given the above information.

Provide a timeline along with your calculations for each step of the question.
Answer #1

a) D1 = 1 x (1 + 2%) = 1.02, D2 = 1.02 x 1.02 = 1.04, D3 = 1.04 x 1.02 = 1.06

b) PV = 1.02 / 1.08 + 1.04 / 1.08^2 + 1.06 / 1.08^3 = $2.68

c) PV = P3 / (1 + r)^3 = 30 / (1 + 8%)^3 = $23.81

d) Price today = Sum of b and c = 2.68 + 23.81 = $26.49